# AFIRE: Accurate and Fast Image Reconstruction Algorithm for Geometric-inconsistency Multispectral CT

### v1

@@ 3.1.2 The proposed algorithm for the continuous problem
Since $\mathcal{P}$ is invertible from $\mathcal{F}$ to $\tilde{\mathcal{G}}$, $\boldsymbol{\mathcal{P}}$ is invertible from $\mathcal{F}^{D}$ to $\tilde{\mathcal{G}}^{Q}$ when $Q=D$. Combining with the invertibility of $\boldsymbol{\varphi}$, we obtain the invertibility of $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$. ∎
@@
The invertibility of $\boldsymbol{\varphi}$ implies $Q=D$, which means that the number of used energy spectra should be equal to the number of basis images. Actually, this necessary condition is commonly satisfied in clinical applications, for example, $Q=D=2$ in DECT.
Moreover, denote $\mathcal{P}^{-1}$ as the inverse of $\mathcal{P}$, and naturally
$\boldsymbol{\mathcal{P}}^{-1}=\operatorname*{\texttt{diag}}(\mathcal{P}^{-1},\ldots,\mathcal{P}^{-1})$
@@
$(\mathcal{D}\boldsymbol{K})^{-1}(\hat{\boldsymbol{f}},\cdot):\tilde{\mathcal{G}}^{Q}$ $\longrightarrow\mathcal{F}^{D}$
@@
Hence, by leveraging the simplified Newton method [10], the proposed iterative scheme for solving 9 can be formulated as
@@
Note that 17 is well-defined if $\boldsymbol{g}-\boldsymbol{K}(\boldsymbol{f}^{k})\in\tilde{\mathcal{G}}^{Q}$ for $\boldsymbol{f}^{k}\in\mathcal{F}^{D}$. Based on the iterative scheme in 17, we present the algorithm AFIRE for solving the continuous MSCT image reconstruction problem in algorithm 1.
@@
1:Initialize: Given the measured data $\{g^{[q]}\}_{q=1}^{Q}$, the energy spectra distributions $\{s^{[q]}\}_{q=1}^{Q}$, the MAC functions $\{b_{d}\}_{d=1}^{D}$ and a certain real constant set $\{C_{d}\}_{d=1}^{D}$. Initialize the basis image $\boldsymbol{f}^{0}$. Let $k\leftarrow 0$.
[…]
@@
$\begin{bmatrix}\varLambda_{1}^{k}\\ \vdots\\ \varLambda_{Q}^{k}\end{bmatrix}\leftarrow\begin{bmatrix}g^{[1]}-K^{[1]}(\boldsymbol{f}^{k})\\ \vdots\\ g^{[Q]}-K^{[Q]}(\boldsymbol{f}^{k})\end{bmatrix}.$
@@
$\begin{bmatrix}f_{1}^{k+1}\\ \vdots\\ f_{D}^{k+1}\end{bmatrix}\leftarrow\begin{bmatrix}f_{1}^{k}\\ \vdots\\ f_{D}^{k}\end{bmatrix}-\boldsymbol{\varphi}^{-1}\begin{bmatrix}\mathcal{P}^{-1}\varLambda_{1}^{k}\\ \vdots\\ \mathcal{P}^{-1}\varLambda_{Q}^{k}\end{bmatrix}.$
@@
In algorithm 1, one can specifically choose the $C_{d}=0$ for $d=1,\ldots,D$ with $\hat{\boldsymbol{f}}=\boldsymbol{0}\in\mathcal{F}^{D}$ to fulfill the condition in 12. The corresponding formula of 14 becomes
$\varphi^{[q]}_{d}=\int_{0}^{E_{\max}}s^{[q]}(E)b_{d}(E)\mathrm{d}E.$
@@ 3.2 The discrete setting
Here, we start from the DD-data model 4 and then present the associated algorithm for solving the geometric-inconsistency full-scan MSCT image reconstruction problem.
@@ 3.2.1 Discrete MSCT image reconstruction problem
As described in section 2.2, for the discrete case, the forward operator of ray $j$ with the $q$-th energy spectrum may be formulated as
@@
$K_{j}^{[q]}(\mathbf{f})=g_{j}^{[q]}\quad\text{for}~{}q=1,\ldots,Q~{}\text{and}~{}j=1,\ldots,J_{q}.$ (19)
The compact form of 19 can be expressed as
@@ 3.2.2 The proposed algorithm for the discrete problem
and $\odot$ denotes the operation of the element-wise product.
@@
Note that probably the most simple example satisfying the condition of 23 is taking $\hat{\mathbf{f}}=\boldsymbol{0}$, which leads to $C_{d}^{[q]}=0$ for any indices $q$ and $d$. But searching for the other points is beyond the scope of this work.
@@
where $\mathbf{P}_{q}=[\mathbf{p}_{1}^{[q]},\dots,\mathbf{p}_{J_{q}}^{[q]}]^{\boldsymbol{\mathsf{T}}}$ is the projection matrix under the $q$-th energy spectrum. Furthermore, the associated Jacobian matrix of $\mathbf{K}$ at $\hat{\mathbf{f}}$ is
@@
Again, by leveraging the simplified Newton method as before, the proposed iterative scheme for solving the discrete problem in 19 (or 20) can be formulated as
@@
1:Initialize: Given the measured data $\{\mathbf{g}^{[q]}\}_{q=1}^{Q}$, the discrete energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$, the MAC $\{\mathbf{b}_{d}\}_{d=1}^{D}$, and a certain real constant set $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$. Initialized the basis images $\{\mathbf{f}_{d}^{0}\}_{d=1}^{D}$. Let $k\leftarrow 0$.
@@
$\begin{bmatrix}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{g}^{[1]}-\mathbf{K}^{[1]}(\mathbf{f}^{k})\\ \vdots\\ \mathbf{g}^{[Q]}-\mathbf{K}^{[Q]}(\mathbf{f}^{k})\end{bmatrix}.$
@@
$\begin{bmatrix}\mathbf{f}_{1}^{k+1}\\ \vdots\\ \mathbf{f}_{D}^{k+1}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{f}_{1}^{k}\\ \vdots\\ \mathbf{f}_{D}^{k}\end{bmatrix}-\boldsymbol{\phi}^{-1}\otimes\begin{bmatrix}\mathbf{P}_{1}^{-1}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \mathbf{P}_{Q}^{-1}\boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}.$
@@
As a matter of fact, the inverse of the matrix $\mathbf{P}$ in 27 is difficult to calculate accurately, as all of its diagonal elements $\mathbf{P}_{1},\ldots,\mathbf{P}_{Q}$ may not even be invertible square matrices, even if they are, the computational cost of calculating the inverse of a large-scale matrix is very high. Therefore, we can only obtain an approximation for the inverse of $\mathbf{P}$, and the actual iterative scheme used in algorithm 2 should be
@@
where $\mathbf{P}^{+}=\operatorname*{\texttt{diag}}(\mathbf{P}_{1}^{+},\ldots,\mathbf{P}_{Q}^{+})$ is the approximation for the inverse of $\mathbf{P}$. Here $\mathbf{P}_{q}^{+}$ denotes the approximation for the inverse of $\mathbf{P}_{q}$. Subsequently, we present the actual algorithm AFIRE used for solving the discrete problem of 19 (or 20) in algorithm 3.
Algorithm 3 The actual algorithm used for solving the discrete problem 19 (or 20).
1:Initialize: Given the measured data $\{\mathbf{g}^{[q]}\}_{q=1}^{Q}$, the discrete energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$, the MAC $\{\mathbf{b}_{d}\}_{d=1}^{D}$, and a certain real constant set $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$. Initialized the basis images $\{\mathbf{f}_{d}^{0}\}_{d=1}^{D}$. Let $k\leftarrow 0$.
3:Loop:
4: Compute the residual
$\begin{bmatrix}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{g}^{[1]}-\mathbf{K}^{[1]}(\mathbf{f}^{k})\\ \vdots\\ \mathbf{g}^{[Q]}-\mathbf{K}^{[Q]}(\mathbf{f}^{k})\end{bmatrix}.$
5: Update $\mathbf{f}^{k+1}$ by
$\begin{bmatrix}\mathbf{f}_{1}^{k+1}\\ \vdots\\ \mathbf{f}_{D}^{k+1}\end{bmatrix}\leftarrow\begin{bmatrix}\mathbf{f}_{1}^{k}\\ \vdots\\ \mathbf{f}_{D}^{k}\end{bmatrix}-\boldsymbol{\phi}^{-1}\otimes\begin{bmatrix}\mathbf{P}_{\!1}^{+}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \mathbf{P}_{\!Q}^{+}\boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}.$
otherwise, let $k\leftarrow k+1$, goto Loop.
@@ 4 Convergence analysis
In this section, we study the convergence of the proposed algorithms that given in algorithms 1, 2 and 3. Assuming that both the solutions exist to the continuous problem 9 and the discrete one 20, and we refer to $\boldsymbol{f}^{*}$ and $\mathbf{f}^{*}$ as the solutions, respectively.
Let $\|\cdot\|_{F}$ and $\sigma_{\min}(\cdot)$ be the Frobenius norm and the minimum singular value of a matrix, respectively. If $\sigma_{\min}(\cdot)$ is nonvanishing, define the scaled condition number by $\kappa_{F}(\cdot):=\|\cdot\|_{F}/\sigma_{\min}(\cdot)$. Denote $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ as the $1$-norm and Euclidean norm of a vector or the spectral norm of a matrix, respectively.
@@ 4.1 Convergence analysis for the continuous setting
Suppose that the conditions in proposition 1 hold and $\boldsymbol{\varphi}$ is invertible. Assume that there exists a constant $\eta>0$ such that, for any nonnegative normalized function $\rho$ and any indices $q$, $d$,
@@
Then, for any $\boldsymbol{f},\boldsymbol{h}\in\mathcal{F}^{D}$,
$\|\mathcal{D}\boldsymbol{K}(\boldsymbol{f},\boldsymbol{h})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathcal{G}^{Q}}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathcal{G}^{Q}}.$ (29)
@@
Suppose that the conditions in lemma 1 hold and $\eta\kappa_{F}(\boldsymbol{\varphi})<1$, and the iterative scheme 17 is well-defined. Then, for any initial point $\boldsymbol{f}^{0}\in\mathcal{F}^{D}$, the iteration sequence $\{\boldsymbol{f}^{k}\}$ generated by algorithm 1 converges to the unique solution of 9.
@@
$\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},$ $\boldsymbol{f}^{k+1}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}$
$=\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k+1}-\boldsymbol{f}^{k})+\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}$
@@
$=\Big{\|}\int_{0}^{1}\mathcal{D}\boldsymbol{K}(t\boldsymbol{f}^{k}+(1-t)\boldsymbol{f}^{*},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\mathrm{d}t\Big{\|}_{\mathcal{G}^{D}}$
$\leq\int_{0}^{1}\|\mathcal{D}\boldsymbol{K}(t\boldsymbol{f}^{k}+(1-t)\boldsymbol{f}^{*},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}\mathrm{d}t$
@@
$\big{\|}\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\big{\|}_{\mathcal{G}^{D}}\longrightarrow 0\quad\text{as}~{}k\longrightarrow\infty.$
@@
$\|\boldsymbol{f}^{k}-\boldsymbol{f}^{*}\|_{\mathcal{F}^{D}}\longrightarrow 0\quad\text{as}~{}k\longrightarrow\infty.$
@@
$\|\boldsymbol{K}(\bar{\boldsymbol{f}}^{*})-\boldsymbol{K}(\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}.$
Thus, $\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathcal{G}^{D}}=0$, which implies that $\bar{\boldsymbol{f}}^{*}=\boldsymbol{f}^{*}$. ∎

### v2

@@ 3.1.2 The proposed algorithm for the continuous problem
Since $\mathcal{P}_{1},\ldots,\mathcal{P}_{Q}$ are invertible, $\boldsymbol{\mathcal{P}}$ is invertible. Combining with the invertibility of $\boldsymbol{\varphi}$, we obtain the invertibility of $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$. ∎
@@
If $\mathbb{S}^{n-1}_{q}\subseteq\mathbb{S}^{n-1}$ meets every equatorial circle of $\mathbb{S}^{n-1}$, it follows from [32, theorem 2.11] that $\mathcal{P}_{q}:\mathcal{F}\rightarrow\tilde{\mathcal{G}}_{q}$ is invertible. The condition of $Q=D$ means that the number of used energy spectra should be equal to the number of basis images, which is commonly satisfied in clinical applications. For example, $Q=D=2$ in DECT.
Moreover, denote $\mathcal{P}^{-1}_{q}$ as the inverse of $\mathcal{P}_{q}$, and naturally
$\boldsymbol{\mathcal{P}}^{-1}=\operatorname*{\texttt{diag}}(\mathcal{P}_{1}^{-1},\ldots,\mathcal{P}_{Q}^{-1})$
@@
$(\mathcal{D}\boldsymbol{K})^{-1}(\hat{\boldsymbol{f}},\cdot):\tilde{\mathbfcal{G}}^{Q}$ $\longrightarrow\mathbfcal{F}^{D}$
@@
Hence, by leveraging the simplified Newton method [12, 33], the proposed iterative scheme for solving 9 can be formulated as
@@
Note that 17 is well-defined if $\boldsymbol{g}-\boldsymbol{K}(\boldsymbol{f}^{k})\in\tilde{\mathbfcal{G}}^{Q}$ for $\boldsymbol{f}^{k}\in\mathbfcal{F}^{D}$, and the standard simplified Newton method considers the derivative at the initial point $\boldsymbol{f}^{0}$. Based on the iterative scheme in 17, we present the algorithm AFIRE for solving the continuous MSCT image reconstruction problem in algorithm 1.
@@
1:Initialize: Given $Q=D$, the measured data $\{g^{[q]}\}_{q=1}^{Q}$, the distributions of energy spectra $\{s^{[q]}\}_{q=1}^{Q}$, the MAC functions $\{b_{d}\}_{d=1}^{D}$ and a certain real constant set $\{C_{d}\}_{d=1}^{D}$. Initialize the basis image $\boldsymbol{f}^{0}$. Let $k\leftarrow 0$.
2:Compute the constant matrix $\boldsymbol{\varphi}=(\varphi^{[q]}_{d})_{Q\times D}$ by 14 and the inverse $\boldsymbol{\varphi}^{-1}$.
@@
$\begin{bmatrix}\varLambda_{1}^{k}\\ \vdots\\ \varLambda_{Q}^{k}\end{bmatrix}\longleftarrow\begin{bmatrix}g^{[1]}-K^{[1]}(\boldsymbol{f}^{k})\\ \vdots\\ g^{[Q]}-K^{[Q]}(\boldsymbol{f}^{k})\end{bmatrix}.$
@@
$\begin{bmatrix}f_{1}^{k+1}\\ \vdots\\ f_{D}^{k+1}\end{bmatrix}\longleftarrow\begin{bmatrix}f_{1}^{k}\\ \vdots\\ f_{D}^{k}\end{bmatrix}-\boldsymbol{\varphi}^{-1}\begin{bmatrix}\mathcal{P}^{-1}_{1}\varLambda_{1}^{k}\\ \vdots\\ \mathcal{P}^{-1}_{Q}\varLambda_{Q}^{k}\end{bmatrix}.$
@@
In algorithm 1, one can specifically choose the $C_{d}^{[q]}=0$ for $d=1,\ldots,D$ and $q=1,\ldots,Q$ with $\hat{\boldsymbol{f}}=\boldsymbol{0}\in\mathbfcal{F}^{D}$ to fulfill the condition of 12. The corresponding formula of 14 becomes
$\varphi^{[q]}_{d}=\int_{0}^{E_{\max}}s^{[q]}(E)b_{d}(E)~\mathrm{d}E.$
@@ 3.2 The discrete setting
Here, we start from the DD-data model 4 and then present the associated algorithm for solving the geometric-inconsistent MSCT image reconstruction problem under the mildly full-scan case.
@@ 3.2.1 Discrete MSCT image reconstruction problem
As described in section 2.2, for the discrete case, the forward operator of ray $j$ under the $q$-th energy spectrum may be formulated as
@@
$K_{j}^{[q]}(\mathbf{f})=g_{j}^{[q]}\quad\text{for}~j=1,\ldots,J_{q}~\text{and}~q=1,\ldots,Q.$ (19)
Here $J_{q}$ is the number of rays under the $q$-th energy spectrum. The compact form of 19 can be expressed as
@@
[…]
Remark 4.
In the geometric-consistent case, the same scanning geometries across different energy spectra allow 20 to be divided into a lot of very small-scale independent subsystems along each X-ray path, enabling high solving efficiency via the two-step DDD method [18]. In contrast, the geometric inconsistency prevents this kind of division, requiring to directly solve a large-scale nonlinear system with $\sum_{q=1}^{Q}J_{q}$ equations (e.g., 294,912 equations considered in section 5.2). If the special structure is ignored, solving such a large-scale nonlinear problem with high precision is quite time-consuming.
@@ 3.2.2 The proposed algorithm for the discrete problem
and $\odot$ denotes the operation of the elementwise product.
@@
Remark 5.
The simplest example that satisfying the condition of 23 is taking $\hat{\mathbf{f}}=\boldsymbol{0}$, which leads to $C_{d}^{[q]}=0$ for any $q$ and $d$. Similarly to remark 1, the explicit identification of other such points is not critical and is beyond the scope of this work.
@@
where $\mathbf{P}_{q}=[\mathbf{p}_{1}^{[q]},\dots,\mathbf{p}_{J_{q}}^{[q]}]^{\boldsymbol{\mathsf{T}}}$ is the projection matrix under the $q$-th energy spectrum. Note that $\mathbf{P}_{1}=\cdots=\mathbf{P}_{Q}$ in the geometric-consistent case, but it fails to hold in the geometric-inconsistent case.
Furthermore, the associated Jacobian matrix of $\mathbf{K}$ at $\hat{\mathbf{f}}$ is
@@
Again, by leveraging the simplified Newton method [12, 33], the proposed iterative scheme for solving the discrete problem in 19 (or 20) can be formulated as
@@
Note that the standard simplified Newton method utilizes the Jacobian matrix at the initial point $\mathbf{f}^{0}$.
@@
1:Initialize: Given $Q=D$, the measured data $\{\mathbf{g}^{[q]}\}_{q=1}^{Q}$, the discrete energy spectra $\{\mathbf{s}^{[q]}\}_{q=1}^{Q}$, the MAC $\{\mathbf{b}_{d}\}_{d=1}^{D}$, and a certain real constant set $\{C_{d}^{[q]}\}_{d,q=1}^{D,Q}$. Initialized the basis images $\{\mathbf{f}_{d}^{0}\}_{d=1}^{D}$. Let $k\leftarrow 0$.
@@
$\begin{bmatrix}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}\longleftarrow\begin{bmatrix}\mathbf{g}^{[1]}-\mathbf{K}^{[1]}(\mathbf{f}^{k})\\ \vdots\\ \mathbf{g}^{[Q]}-\mathbf{K}^{[Q]}(\mathbf{f}^{k})\end{bmatrix}.$
@@
$\begin{bmatrix}\mathbf{f}_{1}^{k+1}\\ \vdots\\ \mathbf{f}_{D}^{k+1}\end{bmatrix}\longleftarrow\begin{bmatrix}\mathbf{f}_{1}^{k}\\ \vdots\\ \mathbf{f}_{D}^{k}\end{bmatrix}-\boldsymbol{\phi}^{-1}\otimes\begin{bmatrix}\mathbf{P}_{1}^{-1}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \mathbf{P}_{Q}^{-1}\boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}.$ (28)
@@
Remark 7.
As a matter of fact, the inverse of the matrix $\mathbf{P}$ in 27 is difficult to calculate accurately. That is because all of its diagonal elements $\mathbf{P}_{1},\ldots,\mathbf{P}_{Q}$ may not even be invertible square matrices, even if they are, the computational cost of calculating the inverse of a large-scale matrix is very high. Therefore, we can only obtain an approximation for the inverse of $\mathbf{P}$, and the actual iterative scheme used for 27 should be
@@
where $\mathbf{P}^{+}=\operatorname*{\texttt{diag}}(\mathbf{P}_{1}^{+},\ldots,\mathbf{P}_{Q}^{+})$ is the approximation for the inverse of $\mathbf{P}$. Here $\mathbf{P}_{q}^{+}$ denotes the approximation for the inverse of $\mathbf{P}_{q}$. The update of 28 should be replaced with
$\begin{bmatrix}\mathbf{f}_{1}^{k+1}\\ \vdots\\ \mathbf{f}_{D}^{k+1}\end{bmatrix}\longleftarrow\begin{bmatrix}\mathbf{f}_{1}^{k}\\ \vdots\\ \mathbf{f}_{D}^{k}\end{bmatrix}-\boldsymbol{\phi}^{-1}\otimes\begin{bmatrix}\mathbf{P}_{1}^{+}\boldsymbol{\Lambda}_{1}^{k}\\ \vdots\\ \mathbf{P}_{Q}^{+}\boldsymbol{\Lambda}_{Q}^{k}\end{bmatrix}.$ (30)
@@ 4 Convergence analysis
In this section, we study the convergence of the proposed algorithms that given in algorithms 1 and 2. Assuming that both the solutions exist to the continuous problem 9 and the discrete one 20, and we refer to $\boldsymbol{f}^{*}$ and $\mathbf{f}^{*}$ as the solutions, respectively.
Let $\|\cdot\|_{F}$ and $\sigma_{\min}(\cdot)$ be the Frobenius norm and the minimum singular value of a matrix, respectively. If $\sigma_{\min}(\cdot)$ is nonvanishing, define the scaled condition number by $\kappa_{F}(\cdot):=\|\cdot\|_{F}/\sigma_{\min}(\cdot)$. Denote $\|\cdot\|_{1}$ and $\|\cdot\|_{2}$ as the 111-norm and Euclidean norm of a vector or the spectral norm of a matrix, respectively.
@@ 4.1 Convergence analysis for the continuous setting
Suppose that the conditions in proposition 2 hold, and $\mathbb{T}^{n}_{q}=\mathbb{T}^{n}$ for $q=1,\ldots,Q$. Assume that there exists a constant $\eta>0$ such that, for any nonnegative normalized function $\rho$ and any indices $q,d$,
@@
Then, for any $\boldsymbol{f},\boldsymbol{h}\in\mathbfcal{F}^{D}$,
$\|\mathcal{D}\boldsymbol{K}(\boldsymbol{f},\boldsymbol{h})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathbfcal{G}^{Q}}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{h})\|_{\mathbfcal{G}^{Q}}.$ (31)
@@
Suppose that the conditions in lemma 1 hold and $\eta\kappa_{F}(\boldsymbol{\varphi})<1$, and the iterative scheme 17 is well-defined. Then, for any initial point $\boldsymbol{f}^{0}\in\mathbfcal{F}^{D}$, the iteration sequence $\{\boldsymbol{f}^{k}\}$ generated by algorithm 1 converges to the unique solution of 9.
@@
$\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},$ $\boldsymbol{f}^{k+1}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}$
$=\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k+1}-\boldsymbol{f}^{k})+\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}$
@@
$=\Big{\|}\int_{0}^{1}\mathcal{D}\boldsymbol{K}(t\boldsymbol{f}^{k}+(1-t)\boldsymbol{f}^{*},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\mathrm{d}t\Big{\|}_{\mathbfcal{G}^{Q}}$
$\leq\int_{0}^{1}\|\mathcal{D}\boldsymbol{K}(t\boldsymbol{f}^{k}+(1-t)\boldsymbol{f}^{*},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}\mathrm{d}t$
@@
$\big{\|}\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\boldsymbol{f}^{k}-\boldsymbol{f}^{*})\big{\|}_{\mathbfcal{G}^{Q}}\longrightarrow 0\quad\text{as}~k\longrightarrow\infty.$
@@
$\|\boldsymbol{f}^{k}-\boldsymbol{f}^{*}\|_{\mathbfcal{F}^{D}}\longrightarrow 0\quad\text{as}~k\longrightarrow\infty.$
@@
$\|\boldsymbol{K}(\bar{\boldsymbol{f}}^{*})-\boldsymbol{K}(\boldsymbol{f}^{*})-\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}\leq\eta\kappa_{F}(\boldsymbol{\varphi})\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}.$
Thus, $\|\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\bar{\boldsymbol{f}}^{*}-\boldsymbol{f}^{*})\|_{\mathbfcal{G}^{Q}}=0$, which implies that $\bar{\boldsymbol{f}}^{*}=\boldsymbol{f}^{*}$. ∎